OpenStudy (anonymous):

s+3/s²+6s-5=L⁻¹ s+3/(s+3)² -14 when completing the square is used, can u show me the next step to determine the L⁻¹?

OpenStudy (sirm3d):

\[\large \mathcal L \left\{ \sinh at \right\}=\frac{ s }{ s^2+a^2 }\] and use the first shifting property of the transform \[\large \mathcal L \left\{ e^{at}f(t) \right\}=F(s-a)\]

OpenStudy (anonymous):

erm...L{cosh at} or L{sinh at}??

OpenStudy (sirm3d):

my mistake. it it cosh at, not sinh at.

OpenStudy (anonymous):

nvm, btw thx a lot for help k, i really appreciate it.. =)

OpenStudy (sirm3d):

\[\large F(s+3)= \frac{ s+3 }{ (s+3)^2-14 }\] so \[\large F(s)=\frac{ s }{ s^2-14 }\]